OpenStudy (anonymous):

Prove that 1+i is a root of the equation z^4+3z^2-6z+10=0

OpenStudy (apoorvk):

substitute (1+i) in place of 'z'. then simplify. you ll see it simplifies to 0.

OpenStudy (anonymous):

I can't see that...?

OpenStudy (mani_jha):

Convert 1+i into the polar form and then substitute. \[\theta=45\] \[1+i=\cos45+isin45\] Use this identity: \[(\cos \theta+i \sin \theta)^{n}=\cos(n \theta)+isin(n \theta)\] Get it?

OpenStudy (apoorvk):

z is any no. complex which is a root of the given expression. now root, as you may remember, means a no. which when substituted in the variable of any expression reduces the expression to zero. so here put in z=1+i so you get: \[(1+i)^4 + 3(1+i)^2 - 6(1+i) +10 = 0\] now see after simplifying whether both the real and imaginary parts add up to zero. you need to solve each term of the equation

OpenStudy (anonymous):

I haven't learned polar form yet..

OpenStudy (anonymous):

I understand apporvk :)

OpenStudy (apoorvk):

now i think using polar form when you have learnt it will actually make it pretty easy. right mani!!

OpenStudy (anonymous):

synthetic division works also